# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v2

@@ A.3 Baseline Network
The baseline function used to compute the learning signal $G_{t}-b(s_{t})$ during training is an MLP with 2 hidden layers of dimension 36. As the return $G_{t}$ is the same for all $t$ by construction, the baseline need only depend on the details of the problem instance, namely $\alpha$ and $\mathcal{C}$. The input to $b(s_{t})$ is a vector composed of $\alpha$ and several statistics of $\mathcal{C}$: the average of the node features $\frac{\sum_{i}\mathbf{x}_{i}}{n}$, the total demand $\sum_{i,j}D_{ij}$, the means and standard deviations of the elements of $D$ and $T$, and the cost component weights $\alpha$ and $1-\alpha$.
A.4 Training Hyper-Parameters
@@
Training was performed using the well-known Adam optimizer Kingma and Ba (2015). Table 7 gives the hyper-parameters used during training. In each epoch of training, we augment the training dataset by applying a set of random transformations. These include rescaling the node positions by a random factor uniformly sampled in the range $[0.4,1.6]$, rescaling the demand magnitudes by a random factor uniformly sampled in the range $[0.8,1.2]$, mirroring node positions about the $y$ axis with probability $0.5$, and revolving the node positions by a random angle in $[0,2\pi)$ about their geometric center.
@@
We also randomly sample $\alpha$ separately for each training city in each epoch. Each time alpha is sampled, it has equal probability of being set to $0.0$, set to $1.0$, and sampled uniformly in the range $[0,1]$. This is to encourage the policy to learn both to handle intermediate cases, and the extreme cases of the passenger perspective and operator perspective.
@@
Number of training epochs 5

### v3

@@ A.3 Value Network
The value function used to compute the advantage $A_{t}=G^{H}_{t}-V(s_{t})$ during training is an MLP with 2 hidden layers of dimension 36. The input to $V(s_{t})$ is a vector composed of $\alpha$ and several statistics of $\mathcal{C}$: the average of the node features $\frac{\sum_{i}\mathbf{x}_{i}}{n}$, the total demand $\sum_{i,j}D_{ij}$, the means and standard deviations of the elements of $D$ and $T$, and the cost component weights $\alpha$ and $1-\alpha$. Along with these, it takes the state vector $\mathbf{s}_{t}$.
We experimented with using an additional neural net head that would compute the value estimate $V(s_{t})$ based on the same node embeddings $Y$ as $\textup{NN}_{ext}$ and $\textup{NN}_{halt}$, but found that this performed slightly worse than using the separate MLP.
Appendix B Training Hyper-Parameters
@@
Training was performed using the well-known Adam optimizer Kingma and Ba (2015). Table 7 gives the hyper-parameters used during training. In each step of training, we augment the training dataset by applying a set of random transformations. These include rescaling the node positions by a random factor uniformly sampled in the range $[0.4,1.6]$, rescaling the demand magnitudes by a random factor uniformly sampled in the range $[0.8,1.2]$, mirroring node positions about the $y$ axis with probability $0.5$, and revolving the node positions by a random angle in $[0,2\pi)$ about their geometric center.
@@
We also randomly sample $\alpha$ separately for each training city in each batch. Each time alpha is sampled, it has equal probability of being set to $0.0$, set to $1.0$, and sampled uniformly in the range $[0,1]$. This is to encourage the policy to learn to handle not only intermediate cases, but also the extreme cases of the passenger perspective and operator perspective.
@@
It is common to incorporate an entropy term in the loss function when training policy gradient models. However, we found that doing so harmed performance of the resulting models on this problem, so we did not use such a term when training the model with PPO. Also note that we do not use generalized advantage estimation as described by Schulman et al. (2017) here, so the parameter $\lambda$ is not needed.
A set of preliminary experiments with a range of values for the discount rate parameter $\gamma$ led us to set this parameter to $\gamma=0.95$.
@@
Number of training iterations 200
Discount rate $\gamma$ 0.95
PPO CLIP threshold $\epsilon$ 0.2
Batch size 256
Horizon 120
Num. epochs 1